OpenStudy (anonymous):

When a ball is thrown, its height in meters "h" after "t" seconds is given by the equation h= vt - 5t^2 where "v" is the initial upwards velocity in meters per second. If v = 7 meters per second, find all values of "t" for which h = 1 meter. Do not round any intermediate steps. Round your answer 2 to decimal places. (If there is more than one answer, enter additional answers with the button that says "or".)

OpenStudy (anonymous):

h=7t-5t² 1=7t-5t² 5t²-7t+1=0 Using quadratic formula: t=(7+sqrt(29))/10 = 1.23851648 s or t=(7-sqrt(29))/10 = 0.161483519 s

OpenStudy (anonymous):

thanxz bro ur a math wiz lol ;-)